OpenStudy (anonymous):

log2 9 + log2 a = log2 11

OpenStudy (anonymous):

\[Log_{2}9+Log _{2}a=\log _{2}11\]

OpenStudy (anonymous):

is that your question?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

\[\log _{2}a=\log _{2}11-\log _{2}9=\log _{2}\frac{ 11 }{ 9 }, a=\frac{ 11 }{ 9 }\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

thanks @satellite73

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

this is what one should expect from you :D keep it up